OpenStudy (anonymous):

Which equation describes the line passing Through the points (-2,0) and (0,4) A: y=2x-2 B: y= -2x - 2 C:y =2x+4 D: Y=. -2x+4

OpenStudy (anonymous):

So what you want to do first is find the slope of the line connecting the two points. This is because in the form y=mx+b, m=slope

OpenStudy (anonymous):

adding more to @Grazes 's comment you can find the slope by using this y2-y1/x2-x1 where y2=4, y1=0,x2=0 and x1=-2

OpenStudy (anonymous):

Erm. Could you find the slope so that we know you're still here?

OpenStudy (anonymous):

answer is D because when put x,y coordinate that equation should be satisfied it means should be zero..when x,y=-2,0 and 0,4 it should be ok.

OpenStudy (anonymous):

\[m=\frac{ y2 - y1 }{ x2 - x1 }=\frac{ 4 - 0 }{ 0 - -2 } = \frac{ 4 }{ 2 } = 2\] \[y - y1 = m(x - x1)\] \[y - 0 = 2(x - -2)\] \[y = 2x +4\]